OpenStudy (anonymous):

The slope of a line is -1/5, and the y-intercept is 5. What is the equation of the line written in general form?

OpenStudy (allank):

Hint: The y-intercept given with full co-ordinates is (0,5)

OpenStudy (lgbasallote):

another hint: slope-intercept form is y = mx + b where m is slope and b is y-intercept

OpenStudy (allank):

Go for it @jewjewbri You know the answer.

OpenStudy (anonymous):

5=-1/5x?

OpenStudy (allank):

Nice try. From the equation y = mx + b, you only need to put in the value of m (slope) and b (y-intercept). The y and x remain as they are.

OpenStudy (anonymous):

@jewjewbri Do you know General form of Linear equation?

OpenStudy (anonymous):

Hmm.... x-5y-5=0?

OpenStudy (anonymous):

Almost right; y = -x/5 + 5 5y = -x + 25 -> x + 5y - 25 = 0

OpenStudy (anonymous):

@jewjewbri Questions, comments ???

OpenStudy (anonymous):

Sorry about that, had leave the com. quick. No comments. But now makes sense. Thank you